OpenStudy (anonymous):

write the first five terms of a1=-5, a2=-1

OpenStudy (cwrw238):

what type of series is this? - if its arithmetic then common difference is 4 and the first 5 are -5, -1,3,7,11

OpenStudy (anonymous):

geometric

OpenStudy (cwrw238):

thats different - the common ratio is -1 / -5 = 1/5 so they are -5 , -1 , -5 * (1/5)^2 , -5 * (1/5)^3 , -5 * (1/5)^4 general form of a GP is a, ar, ar^2, ar^3 .....

OpenStudy (anonymous):

how do you get -5 and -1

OpenStudy (cwrw238):

- you gave them as the first two terms a1 = -5 , a2 = -1

OpenStudy (anonymous):

then I get 1/5 how do I get th eothers after that

OpenStudy (cwrw238):

-5 * 1/5 = -1 eat term is formed by multiplying previous one by 1/5

OpenStudy (anonymous):

ohhh ok i got it so 1/25 is next

OpenStudy (cwrw238):

so 3rd term = a1 * r^2 = -5 *(1/5)^2 = -5 * 1/25 = -1/5

OpenStudy (anonymous):

hay could you help me solve one more i would appreciate it

OpenStudy (cwrw238):

haven't got much time

OpenStudy (anonymous):

find a8 for each arithmatic sequence a1=6 d=2 just what do I need to do

OpenStudy (cwrw238):

ok a8 = a1 + 7d

OpenStudy (anonymous):

and you get the answer

OpenStudy (anonymous):

from that

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

just plug in a1=6 and d = 2

OpenStudy (anonymous):

ok thanks appreciate it